OpenStudy (anonymous):

Two urns contain white balls and yellow balls. The first urn contains 3 white balls and 6 yellow balls and the second urn contains 3 white balls and 8 yellow balls. A ball is drawn at random from each urn. What is the probability that both balls are white?

OpenStudy (anonymous):

1st urn 2nd urn white 3 3 yellow 6 8

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

whats the probability for getting white ball from 1st urn ?

OpenStudy (anonymous):

3/9

ganeshie8 (ganeshie8):

yes, save it. whats the probability for getting white ball from 2nd urn ?

OpenStudy (anonymous):

3/11

ganeshie8 (ganeshie8):

yes ! \(P(W_1 ~ AND~W_2 ) = P(W_1) \times P(W_2)\) since the events are independent

ganeshie8 (ganeshie8):

simply multiply both the probabilities

ganeshie8 (ganeshie8):

\[\dfrac{3}{9} \times \dfrac{3}{11}\]

OpenStudy (anonymous):

.09

ganeshie8 (ganeshie8):

\(\dfrac{1}{11} \approx 0.09\) is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

thanks